A compact car can climb a hill in 10 s. The top of the hill is 30 m higher than the bottom, and the car’s mass is 1,000 kg What is the power output of the car?


1
Expert's answer
2021-07-13T11:37:51-0400
"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{1000\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot30\ m}{10\ s}=29400\ W=29.4\ kW."